OpenStudy (anonymous):

Simplify 6 - 4[x + 1 - (y + 3)] + 5y

OpenStudy (anonymous):

please show step by step

pooja195 (pooja195):

Distribute: =6+(−4)(x)+(−4)(1)+(−4)(−y)+(−4)(−3)+5y =6+−4x+−4+4y+12+5y Combine Like Terms: =6+−4x+−4+4y+12+5y =(−4x)+(4y+5y)+(6+−4+12) Finish it off :)

OpenStudy (anonymous):

thank you!!

pooja195 (pooja195):

:) Welcome To Openstudy :D